# Generating Probabilistic Scenario Programs from Natural Language

### v2

@@ 3.4 Scenic Tools
Each of the program sub parts are fed into the Scenic compiler to check for syntactic correctness. If a subpart fails, a zero-shot prompt with the compiler error message and prior context is passed to the LLM to correct the error up to a certain number of retries. Note that because Scenic is built on top of python, there is no type checking or reference checking during compilation. Once all parts pass, they are stitched together into a final program. This final program is then executed up to a certain number of tries where again a zero-shot prompt is used to explain any execution errors to the LLM. An interesting future direction would be to instead use the abstract syntax tree of the program.
@@ 3.5 Output: CARLA Simulation Execution
For all successful end-to-end runs, ScenarioNL can generate multiple simulations by sampling a single fixed program from the Scenic program. The resulting program is then run in a simulator (in our case CARLA). Each simulation is then rendered with realistic physics, dynamics, and graphics. This allows for further semantic validation of the ScenarioNL output. Our initial experiments with GPT-4 Vision and LLaVA, indicate that a strong Vision-Language Model (VLM) such as GPT-4 can serve as an adequate vision critic. However, we do not include the vision critic in our final system as we have yet to determine where in the system the feedback should be provided (ex. directly to the Scenic program or prior during the ToT reasoning). This is left as future work. Still, with the programs that are generated, downstream applications include tools such as VerifAI (Dreossi et al., 2019) to perform formal analysis of their learning based components to better understand their system’s failure modalities.
@@ 4 Evaluation
4.1 Baseline Evaluations
ScenarioNL was evaluated for all supported configurations of language models and prompting techniques. The results are indicated in Table 1. You can find further details about each prompting technique in the Appendix. The evaluation process incorporated the qualitative and quantitative metrics:
@@ 4.1 Baseline Evaluations
Note that the qualitative metrics Accuracy, Relevance, and Expressiveness (ARE) rely on human evaluations of program outputs for each approach involving Scenic programming experts. They are measured out of five where higher is better. Since all three qualitative factors are imperative in generating useful Scenic code, we propose the ARE score as a method of fairly weighting all three of these factors when evaluating our Scenic programs.
Table 1: Baseline evaluation of simple prompting techniques. No compiler or execution feedback was used to improve outputs further. Inference time is measured in seconds per output. For techniques in which more than one model was evaluated, we average their results as there was no statistical significance between them. On the last row is our Compound AI System, ScenarioNL.
@@
DSPy 0 - - - - -
ScenarioNL 90 4.1 4.5 4.2 4.3 $\pm$ 0.1 109.860
@@ 4.2 Compositional Prompting Evaluation
Given the classification of input descriptions in 3 we perform partial human evaluations only on the easy dataset. While parts of the medium are possible, most are not due to the lack of compatible maps in CARLA. In the hard dataset, we have observed that additional prompting is needed to breakdown advanced driving maneuvers (ex. an overtake can be described as accelerate, switch lane to the left, pass the other vehicle, switch lane to the right, and decelerate). This is required as Scenic does not natively support such maneuvers and instead only supports primitive maneuvers.
@@ 2. Bicyclist
Given the relevant objects, we can then continue with object by object prompts to find the closest matching objects supported in the chosen simulator (CARLA). Observe that by using ToT, we have limited two possible objects which could have been identified as relevant: the rear fender of the Cruise AV which is in fact part of the Cruise AV, and the intersection which is not an object but rather an environmental property which will be later answered in a different question. We repeat the ToT prompting for static objects and environment details.
One of the main features of Scenic is the ability to quantify uncertainty through distributions. Hence, we ask our experts about things that are unknown from the description. Following each unknown quantity or property, we further ask for possible distributions over those values. An example final answer of unknown properties is:
The details missing from the description include weather, time of day,
traffic conditions, road surface conditions, layout of the intersection,
presence of road markings or signs, and the behavior of the Cruise AV
and the bicyclist leading up to the accident. These details are
important in understanding the environment and the sequence of events
that led to the accident.
@@ 5 Conclusion and Future Work
In this work, we propose ScenarioNL, a Compound AI System for generating probabilistic Scenic programs from California DMV crash reports. We choose to represent these scenarios as Scenic programs because of ability to model uncertainty as probability distributions in a probabilistic programming language such as Scenic. Additionally, Scenic includes primitives to represent a wide variety of Cyber Physical Systems and their environments. We demonstrate that simple prompting techniques are not enough for state of the art LLMs to produce correct code in low-resource languages such as Scenic. Instead, we build a system that leverages several prompting strategies, and existing tools, to effectively generate programs in a low-resource setting. Our techniques are not unique to Scenic and can be used for other DSLs.
Our work offers several promising future directions. ScenarioNL supports the generation and rendering of realistic CARLA simulations from traffic accident reports, which will enable more efficient human-in-the-loop or automated end-to-end validation of autonomous driving systems. Exploring how to leverage these simulations in the training of reinforcement-learning based agents is one such direction. Other directions include updating LMQL and Scenic so that we can support more LLMs and more behaviors, respectively.
@@ A.1 Prompting Strategies
Zero-shot The zero-shot approach involves a brief description of the Scenic programming language adapted from the Scenic documentation. Though the documentation includes several method names and descriptions, the approach does not provide any well-formed sample Scenic programs as part of the input.
Few-shot The baseline few-shot approach involves the same brief description of the Scenic programming language adapted from the documentation. In addition to providing the same Scenic tutorial, the prompting provides several hand-picked examples of Scenic programs written for the CARLA simulator based on the CARLA Autonomous Driving Challenge. The structure and number of few-shot examples provided is consistent across all few-shot prompting techniques in order to preserve consistency across all reported statistics.
Retrieval Augmented Generation (RAG) RAG is a technique that enhances the generation process by adding information from relevant documents within a specialized database as additional context. This augmented few-shot approach adapts the same structure as the baseline few-shot approach. However, instead of using the same few-shot example Scenic programs for every input, the program selects the most semantically similar examples from a vector database containing CARLA Challenge Scenic programs. To evaluate semantic similarity, embeddings (calculated using the all-mpnet-base-v2 model (Song et al., 2020; HuggingFace, 2023)) of the example Scenic programs are stored in a vector database (we use Pinecone) and are compared against individual input prompt embeddings using vector dot product as the distance metric.
Hypothetical Document Embeddings (HyDE) Hypothetical Document Embeddings (HyDE) (Gao et al., 2022) intend to improve upon the performance of traditional RAG systems, by converting the model input into the same distribution as the set of data contained in the vector database. This is achieved through an initial generation of a Scenic output using the few-shot technique, followed by performing RAG on the embedding of the baseline few-shot Scenic output instead of the direct natural language input.
Function Calling Prior work in other DSLs demonstrates the improved performance of having a language model write DSL code using a custom code-generation API. A function calling approach API is written in a general-purpose programming language contained within the training distribution of the language model, generating code in the desired DSL output language. More specifically, the OpenAI documentation and other sources (Phan et al., 2021) recommend a function-calling approach that requests the language model to generate output in a language contained within the domain.
@@
Constrained Decoding Constrained decoding involves the specification of constraints that impose significant positive or negative weights on token generation. As a result, we can more strictly ensure specific tokens appear or do not appear, reducing the risks of hallucination or syntax errors. Implementing such restrictions within a programming language is challenging, as variable and function definitions can take any form, potentially bypassing the specified constraints. Our method circumvents this issue by splitting up program generation into different sections, each with a set of structural and token-by-token constraints. This ensures compliance with constraints while maintaining the flexibility needed for creating more diverse programs. Still, we cannot overly constrain the outputs within each section without falling into a templating based strategy such as that by Wang et al. (2023b) or Deng et al. (2023).
@@
DSPy The DSPy framework converts the problem of prompt engineering into a learning problem, to improve the efficiency and performance of the overall process. By framing the task as a learning problem, DSPy is able to optimize the best natural language prompt for a given task. Via this process, DSPy can implicitly leverage many of the techniques described above and several others. In order to do so, DSPy requires a measurable loss function. Without a powerful way of measuring semantic meaning and syntactic correctness between programs, we use word for word matching as our loss function. An interesting future direction could be to explore new loss functions or retry this approach now that ScenarioNL has generated many programs.
A.2 ARE Score Rubric

### v3

@@ 3.4 Scenic Tools
Each of the program sub parts are fed into the Scenic compiler to check for syntactic correctness. If a subpart fails, a zero-shot prompt with the compiler error message and prior context is passed to the LLM to correct the error up to a certain number of retries. Note that because Scenic is built on top of python, there is no type checking or reference checking during compilation. Once all parts pass, they are stitched together into a final program. This final program is then executed up to a certain number of tries where again a zero-shot prompt is used to explain any execution errors to the LLM. An interesting future direction would be to instead use the abstract syntax tree of the program since one can represent a syntactically correct program through a JSON schema thus paving way for a variety of Constrained Decoding tools apart from LMQL, including direct input into OpenAI’s API.
@@ 3.5 Output: CARLA Simulation Execution
For all successful end-to-end runs, ScenicNL can generate multiple simulations by sampling a single fixed program from the Scenic program. Each sample yields concrete values which are drawn from the distributions defined in the programs. This is then run in a simulator (in our case CARLA, however, Scenic does support other simulators). Each simulation is then rendered with realistic physics, dynamics, and graphics. This allows for further semantic validation of the ScenicNL output. Since our output is a concise program, asking ”what-if” questions is as simple as modifying a small portion of the program.
Our initial experiments with GPT-4 Vision and LLaVA, indicate that a strong Vision-Language Model (VLM) such as GPT-4 can serve as an adequate vision critic. However, we do not include the vision critic in our final system as we have yet to determine where in the system the feedback should be provided (i.e. directly to the Scenic program or prior in the natural language space during the ToT reasoning). This is left as future work.
With the programs that are generated, there are numerous downstream applications including using tools such as VerifAI (Dreossi et al., 2019) to perform formal analysis of their learning based components to better understand their system’s failure modalities, learning Operational Design Domains Lee et al. (2020); Torfah et al. (2023) to predict when an accident is imminent, retraining of Reinforcement Learning policies on the new data generated by these programs, and many others.
@@ 4 Evaluation
Baseline Evaluations ScenicNL was evaluated for all supported configurations of language models and prompting techniques. The results are indicated in Table 1. In addition to evaluating widely used prompting techniques, we also selected DSPy Khattab et al. (2023) as an example of an automatic prompt-optimization system. We utilized the BootstrapFewShot and BootstrapFinetune teleprompts with answer_exact_match as a metric or loss function. As it’s highly unlikely a full program can be exactly matched, we split the program into five parts (parameter defintions, scene setup, behaviors, placement, and post conditions) and apply the loss function to each part. This loss function, even by parts, leaves much to be desired and an interesting future direction would be to create a new one that can account for both semantic and syntactic correctness. With Constrained Decoding, we use LMQL and write constraints as Python functions. With Function Calling, we define our own API as an intermediary way to build a Scenic program. You can find further details about each prompting technique in the Appendix A.1.
The evaluation process incorporated the following qualitative and quantitative metrics:
@@
Note that the qualitative metrics of Accuracy, Relevance, and Expressiveness rely on human experts familiar with Scenic to evaluation the program outputs. They are measured out of five and the rubric is provided in the Appendix A.3. Since all three qualitative factors are imperative in generating useful Scenic code, we propose the ARE score as a method of fairly weighting all of these factors when evaluating our Scenic programs.
Table 1: Baseline evaluation of simple prompting techniques. No compiler or execution feedback was used to improve outputs further. Inference time is measured in seconds per output. For techniques in which more than one model was evaluated, we average their results as there was no statistical significance between them. With DSPy, the BootstrapFewShot and BootstrapFinetune methods were used with exact string match as the loss function.
@@
DSPy - Exact match 0 - - - - -
ScenicNL 90 4.1 4.5 4.2 4.3 $\pm$ 0.1 109.860
@@ 4.1 Compositional Prompting Evaluation
Given the classification of input descriptions in 3 we perform partial human evaluations only on the easy dataset. While parts of the medium are possible, most are not due to the lack of compatible maps in CARLA. In the hard dataset, we have observed that additional prompting is needed to breakdown advanced driving maneuvers (e.g. an overtake can be described as accelerate, switch lane to the left, pass the other vehicle, switch lane to the right, and decelerate). This is required as Scenic does not natively support such maneuvers and instead only supports primitive maneuvers. An alternative future direction could be to extend the behaviors of Scenic to support such high-level driving behaviors.
@@ 2. Bicyclist
Given the relevant objects, we then continue with object by object prompts to find the closest matching objects supported in the chosen simulator (CARLA). Observe that by using ToT, we have limited two possible objects which could have been identified as relevant: the rear fender of the Cruise AV which is in fact part of the Cruise AV, and the intersection which is not an object but rather an environmental property which will be later answered in a different question. We repeat the ToT prompting for static objects and environment details.
One of the main features of Scenic is the ability to quantify uncertainty through distributions. Thus, we ask our experts about things that are unknown from the description. Following each unknown quantity or property, we further ask for possible distributions over those values. An example final answer of unknown properties is:
The details missing from the description include weather, time of day, traffic
conditions, road surface conditions, layout of the intersection, presence of
road markings or signs, and the behavior of the Cruise AV and the bicyclist leading
up to the accident. These details are important in understanding the environment
and the sequence of events that led to the accident.
@@ 5 Conclusion and Future Work
We propose ScenicNL, a Compound AI System that generates Scenic programs from California DMV crash reports. We choose to represent these scenarios as Scenic programs because of the ability to model uncertainty as probability distributions in a probabilistic programming language such as Scenic. Additionally, Scenic includes primitives to represent a wide variety of Cyber Physical Systems (CPS) and their environments. We demonstrate that simple prompting techniques are not enough for state-of-the-art LLMs to produce correct code in low-resource settings like Scenic. We build a system that leverages several prompting strategies, and existing tools, to effectively generate programs in this low-resource setting. Our techniques are not unique to Scenic and can be used for other DSLs.
Our work offers several promising future directions. ScenicNL supports the generation and rendering of realistic CARLA simulations from traffic accident reports, which will enable more efficient human-in-the-loop or automated end-to-end validation of autonomous driving systems. Exploring how to leverage these simulations in the training of reinforcement-learning based agents is one such direction. Other directions include updating LMQL or replacing it with a more expressive and performant constrained decoding library such as AICI Moskal et al. (2024), and updating Scenic so that we can support more behaviors. Finally, as Scenic supports other CPS, we can extend this work to support other domains like Robotics and Aviation.
@@ A.1 Prompting Strategies
A.1.1 Zero-shot
The zero-shot approach involves a brief description of the Scenic programming language adapted from the Scenic documentation. Though the documentation includes several method names and descriptions, the approach does not provide any well-formed sample Scenic programs as part of the input.
A.1.2 Few-shot
The baseline few-shot approach involves the same brief description of the Scenic programming language adapted from the documentation. In addition to providing the same Scenic tutorial, the prompting provides several hand-picked examples of Scenic programs written for the CARLA simulator based on the CARLA Autonomous Driving Challenge. The structure and number of few-shot examples provided is consistent across all few-shot prompting techniques in order to preserve consistency across all reported statistics.
A.1.3 Retrieval Augmented Generation (RAG)
RAG is a technique that enhances the generation process by adding information from relevant documents within a specialized database as additional context. This augmented few-shot approach adapts the same structure as the baseline few-shot approach. However, instead of using the same few-shot example Scenic programs for every input, the program selects the most semantically similar examples from a vector database containing CARLA Challenge Scenic programs. To evaluate semantic similarity, embeddings (calculated using the all-mpnet-base-v2 model (Song et al., 2020; HuggingFace, 2023)) of the example Scenic programs are stored in a vector database (we use Pinecone) and are compared against individual input prompt embeddings using vector dot product as the distance metric.
A.1.4 Hypothetical Document Embeddings (HyDE)
Hypothetical Document Embeddings (HyDE) (Gao et al., 2022) intend to improve upon the performance of traditional RAG systems, by converting the model input into the same distribution as the set of data contained in the vector database. This is achieved through an initial generation of a Scenic output using the few-shot technique, followed by performing RAG on the embedding of the baseline few-shot Scenic output instead of the direct natural language input.
A.1.5 Function Calling
Prior work in other DSLs demonstrates the improved performance of having a language model write DSL code using a custom code-generation API. A function calling approach API is written in a general-purpose programming language contained within the training distribution of the language model, generating code in the desired DSL output language. More specifically, the OpenAI documentation and other sources (Phan et al., 2021) recommend a function-calling approach that requests the language model to generate output in a language contained within the domain.
@@ A.1.6 Constrained Decoding
Constrained decoding involves the specification of constraints that impose significant positive or negative weights on token generation. As a result, we can more strictly ensure specific tokens appear or do not appear, reducing the risks of hallucination or syntax errors. Implementing such restrictions within a programming language is challenging, as variable and function definitions can take any form, potentially bypassing the specified constraints. Our method circumvents this issue by splitting up program generation into different sections, each with a set of structural and token-by-token constraints. This ensures compliance with constraints while maintaining the flexibility needed for creating more diverse programs. Specifically, we implement this using LMQL Beurer-Kellner et al. (2023) as it allows us to describe constraints through Python code as opposed to having to directly manipulate per token biases. Still, we cannot overly constrain the outputs within each section without falling into a templating based strategy such as that by Wang et al. (2023b) or Deng et al. (2023).
A.1.7 DSPy
@@
The DSPy framework converts the problem of prompt engineering into a learning problem, to improve the efficiency and performance of the overall process. By framing the task as a learning problem, DSPy is able to optimize the best natural language prompt for a given task. Via this process, DSPy can implicitly leverage many of the techniques described above and several others. In order to do so, DSPy requires a measurable loss function. The ideal loss function would be able to measure the semantic meaning between a program and a natural language description, the program’s syntactic correctness, and our proposed ARE metrics. Designing such a loss function would be an interesting future direction. Presently, we use the default answer_exact_match which checks for an exact string match. As it’s unlikely a correct program can be proposed via an exact string match, we instead breakdown the program into parts and apply the loss function accordingly. Thus, we define the following DSPy Signature:
1class GenScenic(dspy.Signature):
2 """
3 Write Scenic probablistic programs given a natural language description
4 of an autonomous vehicle crash report.
5 """
6 question = dspy.InputField()
7 param_definitions = dspy.OutputField(desc=’Necessary imports and parameter definitions’)
8 scene_setup = dspy.OutputField(desc=’Objects that should be placed in the scene’)
9 behaviors = dspy.OutputField(desc=’Dynamics of the objects in the scene’)
10 placement = dspy.OutputField(desc=’Placement of the dynamic objects in the scene’)
11 post_conditions = dspy.OutputField(desc=’Requires and Terminates clauses’)
Listing 1: DSPy Scenic Program Signature
A.2 Prompt Details
As mentioned in Figure 3, we utilize the composition of Tree-of-Thought Yao et al. (2023) with Role Playing, Zero-Shot Prompting, Few-Shot Prompting, and Constrained Decoding. For the blue rectangles which utilize Tree-of-Thought with a single example, the following meta-prompt is used:
Scenic is a probabilistic programming language for modeling the environments of
autonomous cars. A Scenic program defines a distribution over scenes,
configurations of physical objects and agents. Scenic can also define
(probabilistic) policies for dynamic agents, allowing modeling scenarios where
agents take actions over time in response to the state of the world. We use CARLA
to render the scenes and simulate the agents.
We are going to play a game. For the following questions, imagine that you are
3 different autonomous driving experts. For every question, each expert must
provide a step-by-step explanation for how they came up with their answer.
After all the experts have answered the question, you will need to provide a final
answer using the best parts of each expert’s explanation. Use the following format:
"EXPERT_1:"
"<expert_1_answer>"
"EXPERT_2:"
"<expert_2_answer>"
"EXPERT_3:"
"<expert_3_answer>"
"FINAL_ANSWER:"
"<final_answer>
Here is one example of a Scenic program:"
[…]
QUESTION <insert_qs_num>:
The numbers of the questions in Figure 3 correspond to how one could think of how they would manually reason about a crash. The exact questions asked in our prompts are as follows (with few-shot examples omitted for brevity, if specified):
  1. 1.
    This question utilizes an additional single shot example to demonstrate how to fill out the reasoning template from the meta-prompt. The scenario and text of the single shot example were synthetically generated and human edited.
    Based on the description, what are the main objects that need to be
    included in the scene? Provide step-by-step reasoning then provide
    your final answer as a numbered list. Be concise in your reasoning (no
    more than 1-2 sentences per object).
  2. 2.
    In this question, we use no examples. We also enforce a token limit.
    Based on the relevant objects selected from the original description, what
    are the spatial relationships between the objects? (e.g. car is in front of
    pedestrian, etc.) Are the objects moving or stationary? Are they visible or
    occluded? You can only use the following terms to describe spatial
    relationships: in front of, behind, left of, right of, facing, ahead of,
    behind, visible, and not visible.
    Each expert and the final answer should be provided in the following format:
    SPATIAL_RELATIONSHIPS:
    <spatial_relationships>
    MOVEMENT:
    <movement>
    VISIBILITY:
    <visibility>
    JUSTIFICATION:
    <Q2_JUSTIFICATION>
    FINAL ANSWER:
    <Q2_FINAL_ANSWER>
  3. 3.
    A token limit is enforced.
    What are the main events that happened in the scene? (e.g. car stopped when
    pedestrian crossed the street, a car was driving in a lane then switched
    lanes then made a left turn, etc.). Describe these events using Linear
    Temporal Logic (LTL).
  4. 4.
    1. (a)
      A token limit is enforced.
      What details about the world and environment are missing from the
      description? (e.g. weather, time of day, etc.)
    2. (b)
      A 2-shot synthetically generated and human edited example is used with a token limit.
      For each of the relevant objects, what details about the objects are
      missing from the description that you would need to ask the author about
      in order to create a more accurate scene? What are the main
      environmental factors that need to be included in the scene? Your
      questions should cover dynamics of objects in motion (e.g. speed),
      distances between every pair of objects, and environmental conditions
      (e.g. weather). Provide your questions as a numbered list, but do not
      ask about personal details of any individuals involved.
  5. 5.
    A single shot synthetically generated and human edited example is used with a token limit.
    Based on the missing object information from the user, provide a reasonable
    probability distribution over the missing values. Answer only the questions
    that are about distance between objects, speed, weather, or time. For
    example, if the time of day is missing but you know that the scene is in the
    morning, you could use a normal distribution with mean 8am and standard
    deviation 1 hour (Normal(8, 1)). If the color of the car is missing, you
    could use a uniform distribution over common car color string names. If the
    car speed is missing, you could use a normal distribution with mean around a
    reasonable speed limit for area of the scene and reasonable standard
    deviation, etc.
  6. 6.
    A token limit and constrained decoding for the correct objects are enforced. Note that these limits are not actually due to Scenic, but rather CARLA.
    A user will provide you with probability distributions for missing
    information in a vehicle crash description. Your task is to interpret the
    probability distributions and express them as a Scenic program.
    Scenic can only support the following distributions so you must pick the
    closest matching distribution. Under no circumstance should you use any of
    the other distributions
    "Range(low, high) - Uniform distribution over the real range [[low, high]]"
    "DiscreteRange(low, high) - Uniform distribution over the discreet integer
    range [[low, high]]"
    "Normal(mean, std) - Normal distribution with mean and standard deviation"
    "TruncatedNormal(mean, stdDev, low, high) - Normal distribution with mean
    and standard deviation truncated to the range [[low, high]]"
    "Uniform(value, ) - Uniform distribution over the list of values
    provided."
    "Discrete([[value: weight,  ]]) - Discrete distribution over the list
    of values provided with the given weights (e.g., [[value: 0.5, value: 0.2,
    value: 0.3]])"
    "For weather, Scenic can only support a Uniform or Discrete distribution over
    the following values:
    [’ClearNoon’, ’CloudyNoon’, ’WetNoon’, ’WetCloudyNoon’, ’SoftRainNoon’,
    ’MidRainyNoon’, ’HardRainNoon’, ’ClearSunset’, ’CloudySunset’, ’WetSunset’,
    ’WetCloudySunset’, ’SoftRainSunset’, ’MidRainSunset’, ’HardRainSunset’,
    ’ClearNight’, ’CloudyNight’, ’WetNight’, ’WetCloudyNight’,
    ’SoftRainNight’, ’MidRainyNight’, ’HardRainNight’ ,’DustStorm’]"
    Based on the distributions and original description, define Scenic
    distributions over the uncertain values. Provide values for the parameters
    to your distributions. You may not use any of the other distributions. If
    you cannot find a distribution that matches the missing information, you
    must choose the closest matching distribution.
  7. 7.
    Constrained decoding for the correct town selection and a token limit are used. For the description portion, we do not use constrained decoding. For brevity, we do not repeat the list of default maps and descriptions of CARLA here as they can be found in our code or in CARLA’s public documentation.
    Based on the original description, pick from the following the best
    matching town. You may not choose any other town. If you cannot find a town
    that matches the original description, you must choose the closest
    matching town. Then after selecting a town, provide a high-level
    description (ignoring road names) of where in the town we should replicate
    the original description. For example, if the original description
    specified a highway such as US-101, provide a description about the
    properties of that highway, such as it is a 4 lane road.
    <INSERT_LIST_OF_CARLA_MAPS_AND_DESCRIPTIONS>
  8. 8.
    A token limit is enforced. For brevity, we omit the full list of supported Scenic behaviors here as they can be found in our code or in Scenic’s public documentation.
    Here is a list of the supported behaviors in Scenic. Based on the relevant
    objects and important events, which behaviors do we need to use to recreate
    the original description? You may select more than one behavior as they are
    composable. If you cannot find a behavior that matches the original
    description, you must choose the closest matching behavior.
    <INSERT_LIST_OF_SCENIC_BEHAVIORS>
  9. 9.
    A single synthetically generated example is provided. A token limit is used. For brevity, we omit the full list of vehicles and objects that CARLA support as they can be found in our code or in CARLA’s public documentation
    A user will provide you with a list of main objects from a description. For
    each of the main objects, find the closest matching models from the list
    below. If there are any objects in the original description that you see a
    match for (e.g. a traffic cone), include them in your answer even if they
    are not listed as a main object. Specify your answer as the string value of
    that model. In your final answer, only respond with python code as plain text
    without code block syntax around it.
A.3 ARE Score Rubric